Assume you are the head of an Economic Development Policy team. Suppose that you create two teams to collect data on two developing countries (Viet Nam and India) that need support and resources to help them grow economically. As the boss of the team, you task the two teams to collect sufficient data to help the team to create the New Human Development Index for the two countries.  Suppose after 6 months of data collection the teams came back with information in Table 1 below.  Using the information provided in Table 1, answer the following questions:

Table 1: Data collected by teams

Economic Data Team 1: Viet Nam Team 2: India      
PPP GNI per capita $5,859 $7,060
Life Expectancy 76.5 68.8
Mean years of schooling 8.2 6.4
Expected years of schooling 12.7 12.3
  1. Calculate the Income index, Life expectancy index, and Education index for Viet Nam and India.
  2. From (i) calculate the New Human Development Index (HDI) for Viet Nam and India.
  3. Based on your answer in part (ii), where does the New HDI Rank Viet Nam and India in terms of human development?
  4. Now that you know where these two countries rank in terms of human development, which of the two countries would you allocate more resources to, and why?

NB: Please make sure to watch the videos on New HDI before answering this question.
Answer #1

The Human Development Index (HDI) is a statistic composite index of life expectancy, education, and per capita income indicators, which are used to rank countries into four tiers of human development .

HDI = (Life Expectancy Index X Education Index X Income Index) 1/3

1. Formula to calculate Life Expectancy Index (LEI) = Life Expectancy of a country -20/ 83.2-20

2. Formula to calculate Education Index = (MYSI X EYSI) 1/2 – 0/ 0.951 - 0

          Formula to calculate Mean Years of Schooling Index (MYSI) = Mean years of schooling - 0/ 13.2 - 0

          Formula to calculate Expected Years of Schooling Index (EYSI) = Expected Years of Schooling – 0/ 20.6 – 0

3. Formula to calculate Income Index = Log (Country's GNIpc) – Log ($163) / Log ($108,211) – Log ($163)

First we will find out the HDI for Vietnam:

Life Expectancy Index of Vietnam = Life expectancy of Vietnam -20/ 83.2-20 = 76.5-20/ 83.2-20 = 0.94

Education Index of Vietnam = (MYSI X EYSI) 1/2 – 0/ 0.951 - 0 ;so we have to find out MYSI and EYSI first

        Mean Years of Schooling Index (MYSI) for Vietnam = Mean years of schooling of Vietnam- 0/ 13.2 - 0

                                                                                   = 8.2-0/13.2-0=0.621

       Expected Years of Schooling Index (EYSI) for Vietnam= Expected Years of Schooling – 0/ 20.6 – 0

                                                                                       = 12.7-0/20.6-0 = 0.617

Education Index of Vietnam =( MYSI X EYSI) 1/2 – 0/ 0.951 - 0 = (0.621x0.617)1/2 – 0/ 0.951 - 0

                                                                                             = .651

Income Index of Vietnam= Log (Country's GNIpc) – Log ($163) / Log ($108,211) – Log ($163)

                                      = Log (5859) – Log ($163) / Log ($108,211) – Log ($163)

                                      =3.768/2.212 - 5.03/2.212 = 1.70/2.27 = 0.748

HDI of Vietnam= (Life Expectancy Index X Education Index X Income Index) 1/3

      = (0.94x0.651x.748)1/3 = 0.770

Now we will move to the Indian Case:

Life Expectancy Index of India = Life expectancy of India -20/ 83.2-20 = 68.8-20/ 83.2-20 = 0.772

Education Index of India= (MYSI X EYSI) 1/2 – 0/ 0.951 - 0 ;so we have to find out MYSI and EYSI first

        Mean Years of Schooling Index (MYSI) for India = Mean years of schooling of India- 0/ 13.2 - 0

                                                                                   = 6.4-0/13.2-0=0.4848

       Expected Years of Schooling Index (EYSI) for India= Expected Years of Schooling – 0/ 20.6 – 0

                                                                                       = 12.3-0/20.6-0 = 0.597

Education Index of India =( MYSI X EYSI) 1/2 – 0/ 0.951 - 0 = (0.4848x0.597)1/2 – 0/ 0.951 - 0

                                                                                             = .5656

Income Index of India= Log (Country's GNIpc) – Log ($163) / Log ($108,211) – Log ($163)

                                      = Log (7060) – Log ($163) / Log ($108,211) – Log ($163)

                                      =3.849/2.212 - 5.03/2.212 = 1.74/2.27 = 0.766

HDI of India= (Life Expectancy Index X Education Index X Income Index) 1/3

      = (0.597x0.5656x.766)1/3 = 0.63758

According to the new ranking system updates, India ranks 130 out of 189 countries while Vietnam ranks 116 out of 189 countries.

As the Head of Economic Development Policy Team, I will take some initiatives to improve the HDI ranking of India as it is far behind Vietnam though the percapita GNI is higher for India, and hence more resources will be allocated towards the health and education sector of India.